Base your answers to questions 23 and 24 on the following information The following information about the bonds of the Rmart Company was found in the pages of the New York Times. Current Yield 9.9 Sales in $1,000 148 Last Change Rmart 92 XS 1013/4 23. If you purchased 10 Rmart bonds outstanding that had a total face value of $10,000, how much cash would you receive for interest from the Rmart Company each year? A. $990.00 B. $980.00 C. $950.00 D. $101.75 24. If each Rmart bond has a face value of $1,000, the change in the market value of each bond from the prior trading day was a A. decrease of $75.00. B. decrease of $20.00 C. decrease of $7.50 D. decrease of $0.75. 25. The Sheeba Company issued bonds with a face value of $1,000 and with a stated rate of interest of 18% per year. The interest payment dates for these bonds are November and May 1. The Remote Company bought eight of these bonds three years ago. If Remote is preparing financial statements on December 31, it will need to make an adjusting entry to record the interest that has accrued on the Sheeba bonds. Which o the following entries would be correct? A. Bond Interest Receivable Bond Interest Income Bond Interest Income Bond Interest Income Bond Interest Income $240 $1,440 480 720 $240 $1,440 $480 $ 720 B. Bond Interest Receivable C. Bond Interest Income D. Cash
Answer #1

Answer 23

Total face value $10,000

Bond Rate 9.5%

Annual Interest = $10,000 x 9.5% =$ 950 i.e Option C

Answer 24

Face value per Bond = $ 1,000

Change = - 3/4 i.e = -0.75 on $100

Change in market value on $1,000 = decrease of $ 1,0 x 0.75= - $ 7.5 i.e Option C

Answer 25

Last payout November 1

Interest accrued for 2 months = 8 bonds x $1,000 x 18% x 2/12 = $240 i.e Option A

Bonds Interest Receivable Dr. $240

Bond Interest Income Cr. $240